OpenStudy (anonymous):

Choose the two equations you would use to solve the absolute value equation below. Then solve the two equations. |x – 4| = 12 A. x + 4 = –12 and x + 4 = 12; {–8, 8} B. x + 4 = –12 and x + 4 = 12; {–8, –8} C. x – 4 = 12 and x – 4 = –12; {16, –8} D. x – 4 = 12 and x – 4 = –12; {–8, –16}

OpenStudy (mathstudent55):

To solve the absolute equation \(|X| = k\), where \(X \) is an expression in \(x\), and \(k\) is a non-negative number, solve the compound equation \(X = k\) or \(X = -k\)

OpenStudy (mathstudent55):

Example: Solve \(|2x + 3| = 9\) Change the absolute value equation into the compound equation \(2x + 3 = 9\) or \(2x + 3 = -9\) Now solve the compound equation \(2x = 6\) or \(2x = -12\) \(x = 3 \) or \(x = -6\)

OpenStudy (anonymous):

ok thanks :)

OpenStudy (mathstudent55):

You're welcome. What answer did you get?

OpenStudy (anonymous):

oh I didn't start the question yet lol but Im starting it now

OpenStudy (mathstudent55):

Ok. Just so you understand, my example actually solves my equation completely. That's not what you're being asked to do. All you need to do is the first step where you change the absolute value equation into a compound equation.

OpenStudy (anonymous):

Ok